OpenStudy (anonymous):

if a^4 +2a^2b^2+b^4=100, what is a^2 + b^2

myininaya (myininaya):

(a^2+b^2)^2=100 so a^2+b^2=plus or minus 10

OpenStudy (anonymous):

lols ^

OpenStudy (anonymous):

how did that get 8 medals

myininaya (myininaya):

8-5 medals lol

OpenStudy (anonymous):

talk about doing maths without thinking !

OpenStudy (anonymous):

can a sum of squares be negative?

myininaya (myininaya):

think imaginary!

OpenStudy (anonymous):

still think u are wrong :|

myininaya (myininaya):

a=0 a^2=0 b=isqrt{10} b^2=10i^2=-10 a^2+b^2=0+(-10)=-10

myininaya (myininaya):

what do you think now elecengineer?

OpenStudy (anonymous):

what about the 2a^2b^2

myininaya (myininaya):

remember (a+b)^2=a^2+2ab+b^2 so (a^2+b^2)^2=a^4+2a^2b^2+b^4